OpenStudy (anonymous):

solve: 2x^2= 7x +15 what side do you move? do you move the 2x^2 to the right or move the 7x+15 to the left? :0

OpenStudy (anonymous):

2x^2=7x+15 =>2x^2-7x-15=0 =>2x^2-10x+3x-15=0 =>2x(x-5)+3(x-5)=0 =>(x-5)(2x+3)=0 so we can write x-5=0 =>x=5 or2x+3=0 =>2x=-3 =>x=-3/2